# M2NO: Multiresolution Operator Learning with Multiwavelet-based Algebraic Multigrid Method

### v2

@@ 3.1 Problem definition
We define the operator $T$ as an integral operator with a kernel $K$, where $K:D\times D\to L^{2}$. The action of $T$ on an element $a\in\mathcal{A}$ is given by the integral equation:
@@
This setup allows $T$ to operate within the framework of Hilbert spaces, leveraging the properties of $L^{2}$ spaces to facilitate analysis and computations involving the operator.
An illustrative application of the theory described is the learning of the solution operator for a parametric partial differential equation (PDE). Let $\mathcal{L}_{a}$ denote a differential operator that varies with a parameter $a\in\mathcal{A}$. Consider the general PDE defined as:
$(\mathcal{L}_{a}u)(x)=f(x),\quad x\in D,$ (8)
where $D$ is a bounded, open subset of $\mathbb{R}^{d}$, $f$ is a function residing in an appropriate function space, and the equation is subject to a boundary condition on $\partial D$. Assuming the PDE is well-posed, which implies there exists a unique solution under reasonable assumptions on the data and parameters, we define the operator $F^{\dagger}:\mathcal{A}\to\mathcal{U}$. This operator maps the parameter $a$ to the solution $u$ of (8), effectively encapsulating the dependency of the solution on the parameter within a functional mapping framework.
3.2 Multiwavelet-based Restriction and Prolongation Operators
Given an orthonormal polynomial (OP) basis for $V_{0}^{k}$ as $\{\phi_{0},\phi_{1},\ldots,\phi_{k-1}\}$ with respect to the measure $\mu_{0}$, we can generate the basis for subsequent spaces $V_{n}^{k}$ for $n>1$ through shift and scale (hence the multi-scale aspect) operations on the original basis:
$\phi_{jl}^{n}=2^{\frac{n}{2}}\phi_{j}(2^{n}x-l),\quad j=0,\ldots,k-1,\quad l=0,\ldots,2^{n}-1,\quad\text{w.r.t.}\quad\mu_{n},$ (9)
where $\mu_{n}$ is derived from the shift and scale transformations of $\mu_{0}$. The filter coefficients construct the scaling function $\phi$ and wavelet $\psi$; in our context, functions $\phi$ and $\psi$ are known, and we utilize them to determine the filter coefficients $(H^{(0)},H^{(1)},G^{(0)},G^{(1)})$, as they transform from a fine scale $n=1$ to a coarser scale $n=0$. A uniform measure $\mu_{0}$ version is discussed in [1], and [12] extends it to arbitrary measures by incorporating correction terms $\sum^{(0)}$ and $\sum^{(1)}$.
In one dimension, the operators $H$ and $G$ are constructed as
$H=\begin{pmatrix}H^{(0)}&H^{(1)}\end{pmatrix},\quad G=\begin{pmatrix}G^{(0)}&G^{(1)}\end{pmatrix},$ (10)
$H=H_{y}\otimes H_{x},\quad G=\begin{pmatrix}G_{y}\otimes H_{x}\\ H_{y}\otimes G_{x}\\ G_{y}\otimes G_{x}\end{pmatrix},$ (11)
3.2.1 Low-Pass Filter based Operator
One intuitive way is using the low-pass filter as a restriction operator for the multigrid. The restriction operator and prolongation operator is define as
$I_{h}^{2h}=H,\quad I_{2h}^{h}=(I_{h}^{2h})^{T}=H^{T}.$ (12)
At each decomposition level, it is predominantly the low-pass filters that encapsulate the approximation of the original matrix as illustrated in each one of the levels of scale in Fig 4. In the majority of cases, this methodology proves functional because the residual error associated with the matrices $B$, $C$, and $D$ is almost negligible. However, this property is not maintained by the multigrid method in certain complex scenarios, such as those involving oscillatory coefficients.
Parametrization. With the explicit computation of the low-pass filter $H$, we take the LPF restriction operator as $I_{h}^{2h}=H$ and the corresponding LPF prolongation operator as $I_{2h}^{h}=H^{T}$. It is noteworthy that the prolongation operator is identical for both spaces $V_{0}$ and $V_{1}$, which correspond to the coarsest and the next finer grids, respectively.
3.2.2 Schur complement based Operator
To circumvent these limitations, we introduce a novel set of restriction and prolongation operators, which are founded upon wavelet transformation and the Schur complement. Applying the wavelet transform to both sides of the equation and using the orthogonality of $\mathcal{W}$, we obtain
$(\mathcal{W}\mathcal{L}_{j}\mathcal{W}^{T})\mathcal{W}u=\mathcal{W}f.$ (13)
Denote ~$\tilde{\mathcal{L}}_{j}=\mathcal{W}\mathcal{L}_{j}\mathcal{W}^{T}$, which is partitioned as
~$\tilde{\mathcal{L}}_{j}=\begin{pmatrix}T_{j}&B_{j}\\ C_{j}&D_{j}\end{pmatrix}.$ (14)
Using the block UDL decomposition of ~$\tilde{\mathcal{L}}_{j}$ and the Schur complement [11, 10], the inverse ~$\tilde{\mathcal{L}}_{j}^{-1}$ is expressed as
$u=\begin{pmatrix}H^{T}-G^{T}D_{j}^{-1}C_{j}&G^{T}\end{pmatrix}\begin{pmatrix}(T_{j}-B_{j}D_{j}^{-1}C_{j})^{-1}&0\\ 0&D_{j}^{-1}\end{pmatrix}\begin{pmatrix}H-B_{j}D_{j}^{-1}G\\ G\end{pmatrix}f.$ (15)
The restriction and prolongation operator is defined as
$I_{2h}^{h}=\sqrt{2}(H^{T}-G^{T}D_{j}^{-1}C_{j}),\quad I_{h}^{2h}=\frac{1}{\sqrt{2}}(H-B_{j}D_{j}^{-1}G).$ (16)
For detailed explanations and derivations, please refer to the Appendix B.4.
Parametrization. The multiwavelet-based restriction operator and prolongation operator are approximated using three single-layered CNNs, denoted by $B$ and $C$ and $D^{-1}$, while $H$ and $G$ are precomputed. For a detailed implementation of the model configuration, see Appendix F.2.
3.3 Multiwavelet-based Algebraic Multigrid Method for Discrete PDEs
Employing a linear Finite Element Method (FEM) discretization with mesh size $h=\frac{1}{d}$, the discretized system is described by:
$A_{j}*u_{j}=f_{j},$ (17)
where $u_{j},f_{j}\in\mathbb{R}^{d\times d}$. In this model, * denotes the convolution operation for a single channel, using a $3\times 3$ kernel $A$ determined by the elliptic operator and specific boundary conditions. The inverse of $A*$ aligns with the discrete Green’s function in a linear FEM context. The V-cycle multigrid approach, which can be represented as a conventional convolutional neural network, effectively solves this equation [16, 14].
  1. 1.
    Initialization: Set the initial function $f_{n}$ as $a_{n}$ and the initial guess $u_{j}^{0}(j=0,\ldots,n)$.
  2. 2.
    $u_{n}^{i}=u_{n}^{i-1}+S_{n}^{i}*(f_{n}-A_{n}*u_{n}^{i-1}),\quad i=1,\ldots,s_{n}.$ (18)
  3. 3.
    Multigrid correction: Transition to a coarser grid, where a coarse grid operator $A_{j-1}$ (for $j=1,\ldots,n$) is defined. The residual $r_{j}=f_{j}-A_{j}*u_{j}^{s_{j}}$ is restricted using a multiwavelet-based restriction operator $I_{h}^{2h}$, defined in (16). The correction equation $A_{j-1}e_{j-1}=r_{j-1}$ is then solved over a few relaxation steps $s_{j-1}$. The result, $e_{j-1}$, is prolonged back to the finer grid using the multiwavelet-based prolongation operator $I_{2h}^{h}$, as defined in (16), to update the approximation:
    $f_{j-1}$ $=I_{h}^{2h}(f_{j}-A_{j}*u_{j}^{s_{j}}),$ (19)
    $u_{j-1}^{i}$ $=u_{j-1}^{i-1}+S_{j-1}^{i}*f_{j-1},\quad i=1,\ldots,s_{j-1},$ (20)
    $\tilde{u}_{j}$ $=u_{j}^{s_{j}}+I_{2h}^{h}u_{j-1}^{s_{j-1}}.$ (21)
  4. 4.
    Post-processing: At this stage, one may choose to further prolongate to a finer level in a process known as super-resolution or continue with post-smoothing steps.
The following theorem establishes the convergence properties of the V-cycle multigrid operator. The detailed proof of this theorem is provided in the Appendix C.
Theorem 1 (Operator Convergence [34]).
Let us denote the V-cycle multigrid operator as $\mathcal{L}_{mg}$. The convergence can be succinctly expressed by the inequality
$\lVert u-\mathcal{L}_{mg}(A*u)\rVert_{A}\leq\left(1-\frac{1}{c}\right)\lVert u\rVert_{A},$ (22)
where $c$ represents a constant that remains unaffected by the discretization parameter $h$.
The above theorem affirms the uniform approximation efficacy of $\mathcal{L}_{mg}$ in relation to the pseudo-inverse of $A*$, associated with the Green’s function for the elliptic operator $L_{n}$. The notation $\lVert u\rVert_{A}^{2}=(u,A*u)_{\mathbb{L}^{2}(\Omega)}$ defines the energy norm within the domain $\Omega$.
Table 1: Performance comparison with baselines on benchmarks. $L_{2}$ loss is recorded. For clarity, the best result is in bold and the second best is underlined. The results presented in this table were obtained at the highest resolution settings.
Model 1D 2D
Advection Burgers Darcy Navier-Stoke Diffusion-Reaction
$\nu=1e-3$ $\beta=4.0$ $\nu=1e-3$ $\nu=1e-5$
UNet 3.44E-05 3.83E-05 1.54E-01 3.86E-03 2.15E-02 1.46E-01
FNO-2D 1.68E-05 1.81E-05 1.95E-02 4.95E-03 1.72E-02 1.64E-01
MGNO / / 5.72E-02 3.39E-03 1.27E-02 1.12E-01

### v3

@@ 3.1. Problem Definition
We define the operator $\mathcal{G}$ as an integral operator with a kernel $\kappa$, where $\kappa:D\times D\to L^{2}$. The action of $\mathcal{G}$ on an element $a\in\mathcal{A}$ is given by the integral equation:
(7) $\mathcal{G}a(x)=\int_{D}{\kappa(x,y)a(y)dy}.$
This setup allows $\mathcal{G}$ to operate within the framework of Hilbert spaces, leveraging the properties of $L^{2}$ spaces to facilitate analysis and computations involving the operator.
An illustrative application of the theory described is the learning of the solution operator for a parametric PDE. Let $\mathcal{L}_{a}$ denote a differential operator that varies with a parameter $a\in\mathcal{A}$. Consider the general PDE defined as:
(8) $(\mathcal{L}_{a}u)(x)=f(x),\quad x\in D,$
where $D$ is a bounded, open subset of $\mathbb{R}^{d}$, $f$ is a function residing in an appropriate function space, and the equation is subject to a boundary condition on $\partial D$. Assuming the PDE is well-posed, which implies there exists a unique solution under reasonable assumptions on the data and parameters, we define the operator $\mathcal{G}^{\dagger}:\mathcal{A}\to\mathcal{U}$. This operator maps the parameter $a$ to the solution $u$ of (8), effectively encapsulating the dependency of the solution on the parameter within a functional mapping framework.
3.2. Multiwavelet-based Restriction and Prolongation Operators
Given an orthonormal polynomial basis for $V_{0}^{k}$ as $\{\phi_{0},\phi_{1},\ldots,\phi_{k-1}\}$ with respect to the measure $\mu_{0}$, we can generate the basis for subsequent spaces $V_{n}^{k}$ for $n>1$ through shift and scale (hence the multi-scale aspect) operations on the original basis:
(9) $\phi_{jl}^{n}=2^{\frac{n}{2}}\phi_{j}(2^{n}x-l),$
where $j=0,\ldots,k-1$, $l=0,\ldots,2^{n}-1$ with respect to $\mu_{n}$, which is derived from the shift and scale transformations of $\mu_{0}$. The filter coefficients construct the scaling function $\phi$ and wavelet $\psi$; in our context, functions $\phi$ and $\psi$ are known, and we utilize them to determine the filter coefficients $(H^{(0)},H^{(1)},G^{(0)},G^{(1)})$, as they transform from a fine scale $n=1$ to a coarser scale $n=0$. A uniform measure $\mu_{0}$ is discussed in (02:mw), and (21:mwt) extends it to arbitrary measures by incorporating correction terms $\sum^{(0)}$ and $\sum^{(1)}$.
In one dimension, the operators $H$ and $G$ are constructed as
(10) $H=\begin{pmatrix}H^{(0)}&H^{(1)}\end{pmatrix},\quad G=\begin{pmatrix}G^{(0)}&G^{(1)}\end{pmatrix}$,
(11) $H=H_{y}\otimes H_{x},\quad G=\begin{pmatrix}G_{y}\otimes H_{x}\\ H_{y}\otimes G_{x}\\ G_{y}\otimes G_{x}\end{pmatrix}$,
Our motivation is using the low-pass filter as a restriction operator for the multigrid. The restriction operator and prolongation operator are defined as
(12) $I_{h}^{2h}=H,$
(13) $I_{2h}^{h}=(I_{h}^{2h})^{T}=H^{T}.$
At each decomposition level, it is predominantly the low-pass filters that encapsulate the approximation of the original matrix as illustrated in each one of the levels of scale in Fig 5. In the majority of cases, this methodology proves functional because the residual error associated with the matrices $B$, $C$, and $D$ is almost negligible.
With the explicit computation of the low-pass filter $H$, we take the restriction operator as $I_{h}^{2h}=H$ and the corresponding prolongation operator as $I_{2h}^{h}=H^{T}$. It is noteworthy that the prolongation operator is identical for both spaces $V_{0}$ and $V_{1}$, which correspond to the coarsest and the next finer grids, respectively.
Table 1. Performance comparison with baselines on benchmarks. $L_{2}$ loss is recorded.
Model Advection Burgers Darcy NS Diff-React ERA5 (Temp)
UNet 3.44E-05 3.83E-05 1.54E-01 2.15E-02 1.46E-01 7.55E-01
FNO 1.68E-05 1.81E-05 1.95E-02 1.72E-02 1.64E-01 4.06E-03
MgNO / / 5.72E-02 1.27E-02 1.12E-01 9.76E-01
MWT 9.91E-06 2.07E-05 1.91E-02 1.19E-02 1.15E-01 5.28E-03
LSM 3.44E-02 1.51E-02 4.69E-02 1.61E-02 1.58E-01 4.16E-03
GNOT 1.21E-02 1.08E-02 2.14E-02 4.57E-02 1.47E-02 5.03E-03
ONO 1.23E-02 1.17E-02 2.08E-02 3.93E-02 1.52E-02 4.13E-03
Transolver 1.01E-02 1.09E-02 1.73E-02 3.65E-02 1.52E-02 4.41E-03
AMG 8.89E-03 1.05E-02 1.73E-02 3.42E-02 1.40E-02 /
M2NO 2.35E-06 7.52E-06 9.64E-03 3.61E-03 1.01E-01 2.77E-03
Error Reduction 78.41% 60.49% 49.56% 69.64% 9.82% 31.69%
3.3. Multiwavelet-based Multigrid Method for Discrete PDEs
M2NO is a surrogate neural operator maps from the linear finite element space of input functions $a\in\mathcal{A}$, to the linear finite element space of output functions $u\in\mathcal{U}$. The mapping is defined as:
(14) $\begin{cases}h^{0}=a\in\mathcal{A},\\ h^{l}(a)=\sigma(\mathcal{L}_{mg}^{l}h^{l-1}(a)+B^{l}h^{l-1}(a)+b^{l}),&l=1:L\\ u=\mathcal{G}_{\theta}(a)=\mathcal{L}_{mg}^{L+1}(h^{L}(a))\in\mathcal{U},\end{cases}$
where $\sigma$ denotes the point-wise GELU activation function, and $\mathcal{L}^{\ell}_{mg}$, $B^{l}$, and $b^{l}$ are parameters of the linear operators at each layer.
For the multigrid correction particularly, we employ a linear Finite Element Method (FEM) discretization with mesh size $h=\frac{1}{d}$, the discretized system is described by:
(15) $A_{j}*u_{j}=f_{j},$
where $u_{j},f_{j}\in\mathbb{R}^{d\times d}$. In this model, * denotes the convolution operation for a single channel, using a $3\times 3$ kernel $A$ determined by the elliptic operator and specific boundary conditions. The inverse of AA* aligns with the discrete Green’s function in a linear FEM context. The V-cycle multigrid approach, which can be represented as a conventional convolutional neural network, effectively solves this equation (23:interp; 19:mgnet).
  1. (1)
    Initialization: Set the initial function $f_{n}$ as $a_{n}$ and the initial guess $u_{n}^{0}(j=0,\ldots,n)$.
  2. (2)
    (16) $u_{n}^{i}=u_{n}^{i-1}+S_{n}^{i}*(f_{n}-A_{n}*u_{n}^{i-1}),\quad i=1,\ldots,s_{n},$
    with residual error defined as $r_{j}=f_{n}-A_{n}*u_{n}^{i-1}$.
  3. (3)
    Multigrid correction: Transition to a coarser grid, where a coarse grid operator $A_{j-1}$ (for $j=1,\ldots,n$) is defined. The residual $r_{j}=f_{j}-A_{j}*u_{j}^{s_{j}}$ is restricted using a multiwavelet-based restriction operator $I_{h}^{2h}$, defined in (12). The correction equation $A_{j-1}e_{j-1}=r_{j-1}$ is solved through several relaxation iterations denoted as $s_{j-1}$. In this equation, $e_{j-1}$ represents the error of the linear system; for ease of notation, we refer to it as $u_{j-1}$ throughout the paper. The result, $u_{j-1}$, is prolonged back to the finer grid using the multiwavelet-based prolongation operator $I_{2h}^{h}$, as defined in (13), to update the approximation:
    $f_{j-1}$ $=I_{h}^{2h}r_{j},$
    $u_{j-1}^{i}$ $=u_{j-1}^{i-1}+S_{j-1}^{i}*(f_{j-1}-A_{j-1}*u_{j-1}^{i-1}),$
    $\tilde{u}_{j}$ $=u_{j}^{s_{j}}+I_{2h}^{h}u_{j-1}^{s_{j-1}},$
    where $i=1,\ldots,s_{j-1}$, is the iterative number in each level.
  4. (4)
    Post-processing: At this stage, one may choose to further prolong to a finer level in a process known as super-resolution or continue with post-smoothing steps.
We then give the convergence of the operators within the framework, detailed proof of this can be found in Appendix C.
Theorem 3.1 (Approximation of Continuous Operators by M2NO).
Let $X=H^{s}(\Omega)$ and $Y=H^{s^{\prime}}(\Omega)$ for some $s,s^{\prime}\geq 1$, and let $\sigma\in C(\mathbb{R})$ be a non-polynomial activation function used in the M2NO architecture. Given any continuous operator $\mathcal{G}^{*}:X\to Y$, a compact set $C\subseteq X$, and a desired accuracy $\epsilon>0$, there exists a neural network configuration within the M2NO framework, specifically a network with $n$ neurons per layer, such that
(17) $\inf_{\mathcal{G}\in\mathcal{N}_{n}}\sup_{u\in C}\|\mathcal{G}^{*}(u)-\mathcal{G}(u)\|_{Y}\leq\epsilon,$
where $\mathcal{N}_{n}$ denotes the set of M2NO.
3.4. Model Implementation
In the M2NO, the neural networks $A$, $S$, $B$, $C$ and $D^{-1}$ are implemented as single-layer convolutional neural networks (CNNs). The orthonormal polynomial (OP) basis utilized is the Legendre (Leg) basis, chosen with both uniform and non-uniform measures $\mu_{0}$, in alignment with the MWT(21:mwt) methodologies.
From a computational perspective, the restriction and prolongation operators depicted in Eq (12)-(13) and Fig 1 specifically target the transformation of multiscale and multiwavelet coefficients. Despite this focus on coefficients, the actual input and output of the model are point-wise function samples, specifically $(a_{i},u_{i})$. To address this, we synthesize the data sequence into conceptual functions $f_{a}=\sum_{i=1}^{N}a_{i}\phi_{ji}^{n}$ and $f_{u}=\sum_{i=1}^{N}u_{i}\phi_{ji}^{n}$, positing them within the space $V_{n}^{k}$ where $n=\log_{2}N$. It is critical to note that these functions, $f_{a}$ and $f_{u}$, are employed primarily for theoretical articulation and not for direct computation, serving as a notational convention rather than functional entities in the model’s operations.
Choosing Multiwavelet Basis. We have experimented with multiple wavelet bases, including Haar wavelets, Daubechies 4-tap wavelets, and multi-wavelets derived from Legendre and Chebyshev polynomials, all of which have demonstrated effectiveness in PDE-related applications. After careful evaluation, we selected Legendre polynomial-based multi-wavelets due to their superior empirical performance and compatibility with our M2NO architecture. The Legendre basis is particularly advantageous because its orthogonality and polynomial approximation properties provide an efficient, compact representation of PDE operators, significantly improving the model’s ability to capture both global and local solution features. Additionally, we offer Chebyshev polynomial-based multi-wavelets as an optional hyperparameter choice, which also consistently outperform previous methods, underscoring the robustness and flexibility of our approach.
4. Experiments
In this section, we evaluate the performance of the M2NO across a range of PDE benchmarks that incorporate various boundary conditions and conduct Multiresolution Analysis. Additionally, we delve into Spectral Analysis to assess M2NO’s effectiveness in managing frequency-based discrepancies and explore its capabilities as a Preconditioner.
(18) $\min_{\theta\in\Theta}\mathcal{L}(\theta):=\min_{\theta\in\Theta}\frac{1}{N}\sum_{i=1}^{N}\left[\lVert\tilde{\mathcal{G}}_{\theta}(a_{i})-u_{i}\rVert^{2}\right].$
Benchmarks. To substantiate the superiority of our method, we conducted comprehensive method comparisons across multiple benchmark scenarios (21:fno; 22:bench). For Advection, Burgers and Darcy datasets, the split is 60% for training, 20% for validation, and 20% for testing. For the Diffusion-Reaction and Navier-Stokes datasets, the data is allocated with 80% used for training, 10% for validation, and 10% for testing. For ERA5 dataset, we use 12 days for training, 4 days for validation and 4 days for testing.
Baselines. We conducted a comprehensive comparison against established methods including UNet (15:unet), Fourier Neural Operator (FNO) (21:fno), Multiwavelet Transform (MWT) (21:mwt), and Multigrid Neural Operator (MgNO) (23:mgno). We also included transformer-based methods Latent Spectral Model (LSM) (23:LSM), Transolver (24:Transolver), GNOT (23:GNOT), ONO (24:ImprovedOL) and AMG (25:HarnessingSP). All baselines were reproduced from their official repositories, and for each we performed a light hyper-parameter sweep in the vicinity of the authors’ published configurations to report the strongest reproducible numbers. Transolver and LSM were originally released only for 2-D grids; following the AMG (25:HarnessingSP) setting, we treat the 1-D domain as an unstructured point cloud so these models can be evaluated on the 1-D benchmarks. To ensure fairness, every model on a given dataset is trained under an identical optimisation protocol: 500 epochs, mean-squared ($L_{2}$) loss (21:fno), ADAM optimiser (14:adam) with initial learning rate $10^{-3}$ decayed by a factor $\gamma=0.5$ every 100 epochs, and batch size 512 (1-D) or 16 (2-D). Hardware settings are also unified: Darcy and ERA5 runs use a single NVIDIA A6000 (48 GB), while all other experiments run on a single NVIDIA RTX 4090 (24 GB). Because of the very high resolution of ERA5, several large Transformer-based baselines could not be trained with their original widths—or in some cases could not finish training—within the available memory budget; reported ERA5 numbers therefore correspond to the largest configuration that fit on the GPU.
Time-dependent tasks. In our benchmarks, all tasks are time-dependent, except for Darcy Flow. These tasks are auto-regressive, requiring sequential predictions for each subsequent time step. Accordingly, we employ the 2D versions of the Multiwavelet-based Multigrid Neural Operator (M2NO), along with the Fourier Neural Operator (FNO) and Multiwavelet Transform (MWT) models, to predict the next time frame in a step-by-step manner across all benchmarks.
4.1. Main Results
Table 1 presents the performance comparison of M2NO against baseline methods on a diverse set of PDE benchmarks, including Advection, Burgers, Darcy, Navier-Stokes (NS), Diffusion-Reaction (Diff-React), and the large-scale climate dataset ERA5. M2NO achieves the lowest $L_{2}$ loss across all tasks, with notable improvements of up to 78.41% on Advection and 69.64% on Navier-Stokes over the next best baseline. For high-resolution datasets like ERA5, M2NO surpasses Transolver, the closest competitor, with a margin of 31.69% in $L_{2}$ loss, demonstrating its robustness in capturing multi-scale features effectively. These results highlight M2NO’s superior ability to manage both low- and high-frequency error components, providing consistent and scalable accuracy across different PDE types and resolutions. Detailed spectral analysis and multi-resolution studies further substantiate these improvements, as shown in the following sections.
Real-World Data. Table 2 presents the quantitative results on the ERA5 test set, across five prediction targets: temperature (L1 and L2 error), wind U, wind V, and velocity magnitude (L2 error). M2NO achieves the lowest error across all metrics. Notably, for temperature prediction, M2NO reduces the L2 error by 31.69% compared to the next best baseline (LSM), and achieves a similar 33.92% improvement in L1 error, highlighting its superiority in capturing both average and extreme values. For wind components, M2NO achieves 17.26% and 14.60% lower L2 error than the best baseline, demonstrating robust modeling of fine-scale atmospheric flows. In velocity magnitude, M2NO shows a 17.18% reduction in error, further illustrating its ability to resolve complex vector fields. These consistent improvements indicate that M2NO is more effective in learning multi-scale dynamics and reconstructing high-resolution physical fields from sparse or noisy data.
Table 2. Performance comparison with baselines on ERA5. We evaluate four physical variables including temperature, wind components (U, V), and velocity magnitude.
Model Temperature Wind U Wind V Velocity
L1 L2
Unet 7.21E-02 8.31E-02 7.60E-02 2.32E-01 7.55E-01
FNO 3.04E-03 4.06E-03 1.13E-01 3.88E-01 7.85E-01
MGNO 9.76E-01 9.76E-01 1.83E-01 2.43E-01 7.79E-01
MWT 3.63E-03 5.28E-03 1.19E-01 3.72E-01 7.87E-01
LSM 2.80E-03 4.16E-03 9.55E-02 2.43E-01 7.55E-01
Transolver 3.06E-03 4.41E-03 9.90E-02 2.42E-01 7.87E-01
M2NO 1.85E-03 2.77E-03 6.29E-02 1.98E-01 6.25E-01
Error Reduction 33.92% 31.69% 17.26% 14.60% 17.18%
Figure 2. Performance showcases on the ERA5 (above) and NS (below).
Showcases. Additional qualitative evaluation is provided in Figure 2. M2NO achieves significantly lower absolute prediction errors compared to baselines on both ERA5 and NS datasets, capturing finer details and complex structures more accurately. The visualizations clearly demonstrate M2NO’s improved capability to model intricate spatial features inherent in multi-scale and high-resolution PDE solutions.
4.2. Multiresolution Analysis
Table 3. L2 errors of different models on the ERA5 across increasing spatial resolutions.
Model $\times32$ $\times64$ $\times128$ $\times256$ $\times512$
UNet 1.33E-01 9.85E-02 8.94E-02 8.54E-02 8.31E-02
FNO 4.28E-03 4.10E-03 4.23E-03 4.14E-03 4.06E-03
MgNO 1.52E-02 5.24E-01 9.01E-01 9.76E-01 9.76E-01
MWT 8.09E-03 5.46E-03 5.05E-03 5.04E-03 5.28E-03
LSM 4.89E-03 3.97E-03 4.42E-03 4.82E-03 4.16E-03
Transolver 4.03E-03 3.95E-03 4.03E-03 3.85E-03 4.41E-03
M2NO 3.73E-03 3.68E-03 3.24E-03 2.88E-03 2.77E-03
Error Reduction 7.59% 6.64% 19.54% 25.11% 31.69%
Performance Across Multiple Resolutions. Table 3 presents the $L_{2}$ errors of different models on the ERA5 dataset under varying spatial resolutions, ranging from $32\times 32$ to $512\times 512$. M2NO consistently achieves the lowest error across all resolutions, demonstrating its robustness and adaptability in capturing complex multi-scale patterns without the need for parameter adjustments. Notably, the improvement margin widens as resolution increases, reaching a significant 31.69% promotion over the best baseline at $512\times 512$.
Table 4. Computational efficiency comparison across models (measured with input size $\times512$$\times$3, batch size 16).
Model Param Count Param (MB) GPU Mem (MiB) Train (s/epoch) Inference (s/epoch)
UNet 31,037,633 118.44 29,738 13.31 1.45
FNO 1,444,017 10.97 15,740 5.36 0.87
MgNO 227,403,798 867.47 43,768 42.73 3.34
MWT 5,089,281 37.47 23,480 20.77 2.16
LSM 4,801,089 18.33 42,930 62.54 2.99
Transolver 279,607 1.06 41,450 25.54 3.14
M2NO (Ours) 417,377 1.59 17,992 18.39 1.78
Table 5. Super-resolution benchmarks on ERA5: $L_{2}$ loss across resolutions trained at $32\times 32$.
Model $\times64$ $\times128$ $\times256$ $\times512$
FNO 1.77E-01 3.20E-01 3.29E-01 3.33E-01
MWT 2.86E-01 3.39E-01 5.00E-01 6.64E-01
LSM 2.74E-02 4.69E-02 6.17E-02 6.84E-02
M2NO 1.72E-02 2.28E-02 2.52E-02 2.63E-02
Error Reduction 37.10% 51.42% 59.14% 61.52%
Super-Resolution Performance. Table 5 presents the super-resolution benchmarks for ERA5, with models trained at a base resolution of $32\times 32$ and tested across higher resolutions. M2NO consistently achieves the lowest $L_{2}$ losses at each upscale step, outperforming FNO, MWT, and LSM by significant margins. Notably, M2NO’s error reduction is more pronounced as the resolution increases, achieving over 60% improvement at $512\times 512$. This highlights M2NO’s robust capacity to recover fine-scale details in super-resolution tasks, benefiting from its multiwavelet-based multigrid architecture to effectively capture both local and global structures.
4.3. Spectral Analysis
We evaluate M2NO’s ability to reduce high-frequency errors compared to other neural operator models using a two-dimensional Fast Fourier Transform (FFT) to analyze the radial energy spectrum (Fig. 3). The results show that while all models struggle with low-frequency components, M2NO consistently achieves lower error energy across both low and high frequencies. This reflects its strength in capturing large-scale structures and resolving fine-scale details, making it particularly effective for precision-demanding tasks like Navier-Stokes simulations.
@@ 4.3. Spectral Analysis
Figure 3. Models’ Error Energy Spectrum on the NS.
4.4. M2NO as Preconditioner
We evaluate the effectiveness of M2NO as a preconditioner within the GMRES solver for solving a 512-dimensional Poisson equation. Specifically, we use the trained M2NO model to approximate the inverse operator by providing an identity matrix $\mathbf{I}$ as input and utilizing the output $\mathcal{G}(\mathbf{I})$ as the preconditioning matrix $\mathbf{M}$ (25:NPO). This learned preconditioner is then integrated into the GMRES solver to improve convergence. Baseline methods, including traditional numerical preconditioners (Gauss-Seidel, Schwarz) and other neural operator models (UNet, FNO, Transolver), are similarly integrated into GMRES for direct comparison.
Figure 4 compares the convergence behaviors across these methods. M2NO achieves residuals below $10^{-11}$ in approximately $50$ iterations, significantly outperforming all baselines. Traditional preconditioners require substantially more iterations to reach comparable residuals, and neural baselines converge more slowly and plateau at higher residual levels. These results highlight M2NO’s capability to provide efficient and accurate preconditioning through its multiwavelet-based multigrid structure, significantly enhancing solver efficiency for complex, high-dimensional PDEs.